# Hehong Chen, Junfeng Tian, Qian Qi, Ji Zhang, Fei Huang DAMO Academy, Alibaba Group

### v2

@@ 4.1 Experimental Setup
[…]
@@ 4.3 Ablation Study
IU Instruction Understanding
1. Understand text instruction.
2. Do not require a correct answer, but the response
    should be related to the instruction.
VU Visual Understanding
1. Identify image information.
2. The answer faithfully reflects over 60% of visual
    information in the image.
OCR Optical Character Recognition
1. Recognize text information in the image.
2. The answer faithfully reflects over 60% of text
   information in the image.
KTA Knowledge Transfer Ability
1. Transfer knowledge between language and vision.
   (1) understand textual and visual content
   (2) align and transfer visual and language knowledge
2. Answers are mostly accurate with accuracy rate over 80%.
RA Reasoning Ability
1. Combine image and text for reasoning.
   (1) understand textual and visual content
   (2) conduct multi-step reasoning
   (3) generate answers based on multi-step reasoning process
2. The final answer is essentially correct, but it lacks an
   explicit reasoning process. Alternatively, the final answer is
   mostly correct, and the reasoning process is over 80% accurate.
3. For example
   (1) Commonsense Knowledge Reasoning
   (2) Counterfactual Reasoning
   (3) Spatial Relation Reasoning
   (4) Numerical Computation
   (5) Coding
MDA Multi-turn Dialogue Ability
1. Understand instructions and handle multi-turn conversations.
2. It includes clear references to multiple conversations and
   handles natural language semantics in context effectively.
3. The semantics and references are mostly correct, with an
   accuracy rate of over 80%.
@@ 3. The semantics and references are mostly correct, with an
r1
$58.5_{{\color[rgb]{1,0,0}(-41.5)}}$
$38.1_{{\color[rgb]{1,0,0}(-57.1)}}$
$13.3_{{\color[rgb]{1,0,0}(-43.4)}}$
$16.7_{{\color[rgb]{1,0,0}(-70.8)}}$
$17.1_{{\color[rgb]{1,0,0}(-62.9)}}$
$40.0_{{\color[rgb]{1,0,0}(-55.0)}}$
r2
$93.9_{{\color[rgb]{1,0,0}(-6.1)}}$
$47.6_{{\color[rgb]{1,0,0}(-47.6)}}$
$23.3_{{\color[rgb]{1,0,0}(-33.4)}}$
$29.2_{{\color[rgb]{1,0,0}(-58.3)}}$
$14.3_{{\color[rgb]{1,0,0}(-65.7)}}$
$45.0_{{\color[rgb]{1,0,0}(-50.0)}}$
r3
$93.0_{{\color[rgb]{1,0,0}(-7.0)}}$
$73.0_{{\color[rgb]{1,0,0}(-22.2)}}$
$40.0_{{\color[rgb]{1,0,0}(-16.7)}}$
$41.7_{{\color[rgb]{1,0,0}(-45.8)}}$
$48.6_{{\color[rgb]{1,0,0}(-31.4)}}$
$80.0_{{\color[rgb]{1,0,0}(-15.0)}}$
r4
$86.6_{{\color[rgb]{1,0,0}(-13.4)}}$
$68.3_{{\color[rgb]{1,0,0}(-26.9)}}$
$40.0_{{\color[rgb]{1,0,0}(-16.7)}}$
$50.0_{{\color[rgb]{1,0,0}(-37.5)}}$
$60.0_{{\color[rgb]{1,0,0}(-20.0)}}$
$75.0_{{\color[rgb]{1,0,0}(-20.0)}}$

### v3

# Hehong Chen, Junfeng Tian, Qi Qian, Ji Zhang, Fei Huang, Jingren Zhou DAMO Academy, Alibaba Group

@@ 4.1 Experimental Setup
  • BLIP-2 (Li et al., 2023) is pre-trained through bootstrapped learning from off-the-shelf frozen pre-trained image models and large language models using an efficient pre-training strategy. We use the released code of BLIP-2 ViT-G FlanT5XXL [^2] https://github.com/salesforce/LAVIS/tree/main/projects/blip2 to perform zero-shot generation.
@@ 4.3 Ablation Study
IU Instruction Understanding 1. Understand text instruction. 2. Do not require a correct answer, but the response       should be related to the instruction.
VU Visual Understanding 1. Identify image information. 2. The answer faithfully reflects over 60% of visual       information in the image.
OCR Optical Character Recognition 1. Recognize text information in the image. 2. The answer faithfully reflects over 60% of text      information in the image.
KTA Knowledge Transfer Ability 1. Transfer knowledge between language and vision.      (1) understand textual and visual content      (2) align and transfer visual and language knowledge 2. Answers are mostly accurate with accuracy rate over 80%.
RA Reasoning Ability 1. Combine image and text for reasoning.      (1) understand textual and visual content      (2) conduct multi-step reasoning      (3) generate answers based on multi-step reasoning process 2. The final answer is essentially correct, but it lacks an      explicit reasoning process. Alternatively, the final answer is      mostly correct, and the reasoning process is over 80% accurate. 3. For example      (1) Commonsense Knowledge Reasoning      (2) Counterfactual Reasoning      (3) Spatial Relation Reasoning      (4) Numerical Computation      (5) Coding
MDA Multi-turn Dialogue Ability 1. Understand instructions and handle multi-turn conversations. 2. It includes clear references to multiple conversations and      handles natural language semantics in context effectively. 3. The semantics and references are mostly correct, with an      accuracy rate of over 80%.
@@
r1 $58.5_{{\color[rgb]{1,0,0}(-41.5)}}$ $38.1_{{\color[rgb]{1,0,0}(-57.1)}}$ $13.3_{{\color[rgb]{1,0,0}(-43.4)}}$ $16.7_{{\color[rgb]{1,0,0}(-70.8)}}$ $17.1_{{\color[rgb]{1,0,0}(-62.9)}}$ $40.0_{{\color[rgb]{1,0,0}(-55.0)}}$
r2 $93.9_{{\color[rgb]{1,0,0}(-6.1)}}$ $47.6_{{\color[rgb]{1,0,0}(-47.6)}}$ $23.3_{{\color[rgb]{1,0,0}(-33.4)}}$ $29.2_{{\color[rgb]{1,0,0}(-58.3)}}$ $14.3_{{\color[rgb]{1,0,0}(-65.7)}}$ $45.0_{{\color[rgb]{1,0,0}(-50.0)}}$
r3 $93.0_{{\color[rgb]{1,0,0}(-7.0)}}$ $73.0_{{\color[rgb]{1,0,0}(-22.2)}}$ $40.0_{{\color[rgb]{1,0,0}(-16.7)}}$ $41.7_{{\color[rgb]{1,0,0}(-45.8)}}$ $48.6_{{\color[rgb]{1,0,0}(-31.4)}}$ $80.0_{{\color[rgb]{1,0,0}(-15.0)}}$
r4 $86.6_{{\color[rgb]{1,0,0}(-13.4)}}$ $68.3_{{\color[rgb]{1,0,0}(-26.9)}}$ $40.0_{{\color[rgb]{1,0,0}(-16.7)}}$ $50.0_{{\color[rgb]{1,0,0}(-37.5)}}$ $60.0_{{\color[rgb]{1,0,0}(-20.0)}}$ $75.0_{{\color[rgb]{1,0,0}(-20.0)}}$